The Vinny Cartier Company issued bonds at $1,000 per bond. The bonds had a 25-year life when issued, with semiannual payments at the then annual rate of 9 percent. This return was in line with required returns by bondholders at that point, as described below: Real rate of return 2 % Inflation premium 4 Risk premium 3 Total return 9 % Assume that ten years later the inflation premium is 3 percent, the risk premium has declined to 2 percent and both are appropriately reflected in the required return (or yield to maturity) of the bonds. The bonds have 15 years remaining until maturity. Compute the new price of the bond. Use Appendix B and Appendix D. (Round "PV Factor" to 3 decimal places. Do not round intermediate calculation. Round the final answer to 2 decimal places.)

New price of the bond $
Answer #1

Value of bond is equal to the present value of all remaining coupon payments and the principal payment

New semi annual required return = (2+3+2)/2 = 3.5%

Number of semi annual periods remaining = 15*2 = 30

Hence, new price = 1000*4.5%*PVAF(3.5%, 30 periods) +1000*PVF(3.5%,30 periods)

=45*18.3920 +1000*0.3563

=$1,183.94